OpenStudy (anonymous):

If f(x)=x^2+1 and g(x)=3x+1, find 2f(1)+3g(4).

OpenStudy (babalooo):

you seem familiar.. didn't get the last question huh

OpenStudy (babalooo):

okay i''m not gonna solve the thing for you... someone else prob will I'll just help with understnding.. you know f(x) and g(x) are expression and when they say f(1) or g(4) that just means to find the solution to the f(x) and g(x) by replacing the x variable in the expressions with the number inside the brackets. so if f(x) = x+1 then f(1)=1+1 =2 so f(1) = 2 and 2f(1) is just 2(2)=4. Hope this help you understand

OpenStudy (anonymous):

Yup

OpenStudy (babalooo):

if you solve your question, I can help you check if you're right or not

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Ok so far f(2) and g(13) is that correct

OpenStudy (babalooo):

you mean f(1) =2 and g(4)=13 right. that is correct. f(2) and g(13) means something totally different

OpenStudy (anonymous):

Oh ok

OpenStudy (babalooo):

Yup, now you just have to sub in into the last equation and you're done

OpenStudy (anonymous):

That's what I don't get that part never could get past that.

OpenStudy (babalooo):

okay so you got f(1)=2 and g(4)=13 right. Let's just make sure we are on the same page.

OpenStudy (babalooo):

and you understand how you got those values right

OpenStudy (anonymous):

Yes

OpenStudy (babalooo):

and we are looking for 2f(1)+3g(4). You may think this is confusing. but try seeing it as 2a+3b where a=f(1)=2 and b=g(4)=13. does it look solvable to you now?

OpenStudy (anonymous):

So I'm just going to add 2+13

OpenStudy (babalooo):

well not exactly. it's 2a+3b. where a=2 for f(1) and b=13 for g(4) so it should be 2(2) + 3(13). see what I did there? do you understand?

OpenStudy (anonymous):

Oh ok

OpenStudy (babalooo):

Let me know if you got the final answer

OpenStudy (anonymous):

43

OpenStudy (babalooo):

Yup!

OpenStudy (aum):

Here is the complete solution in one place: f(x) = x^2 + 1 f(1) = 1^2 + 1 = 2 g(x) = 3x + 1 g(4) = 3(4) + 1 = 13 Plug f(1) and g(4) into 2f(1) + 3g(4): 2f(1) + 3g(4) = 2(2) + 3(13) = 4 + 39 = 43

OpenStudy (anonymous):

Thanks a lot